State whether the following statements are true or false. Justify your answer with the help of a short proof or a counter example. i) Equation cos(x+y)p + sin(x+y)q = z^2 is a quasi-linear equation. ii) The solution of PDE dz/dx + dz/dy =z^2 is z = -[y+f(x-y)] Iii)(dz/dx)(dz/dy) - (dz/dy)^2 = 0 is a non-linear PDE
Please do it stepwise as I am a noob in differential equations
1
Expert's answer
2020-04-13T11:02:41-0400

QUESTION(i)

Recall the theory.

Definition. Quasi-linear partial differential equations of first order can be written as



"P(x,y,z)p+Q(x,y,z)q=R(x,y,z)\quad\quad (1)"

where "P,Q,R" are functions of "x,y" and "z" . They do not involve "p" or "q" . Here, linear means that "p" and "q" appear to the first degree only.

Definition taken from book : E. Rukmangadachari. Mathematical Methods


In our case,



"\boxed{\cos(x+y)p+\sin(x+y)q=z^2-\text{s a quasi-linear equation}}"

QUESTION(ii)



"z=-y-f(x-y)\longrightarrow\\[0.3cm]
\left\{\begin{array}{l}
u=x-y\\[0.3cm]
\displaystyle\frac{\partial z}{\partial x}=-\displaystyle\frac{\partial f}{\partial u}\cdot
\displaystyle\frac{\partial u}{\partial x}\equiv-f' \\[0.3cm]
\displaystyle\frac{\partial z}{\partial y}=-1-\displaystyle\frac{\partial f}{\partial y}\cdot
\displaystyle\frac{\partial u}{\partial y}\equiv-1-f'\cdot(-1)
\end{array}\right.\longrightarrow\\[0.3cm]
\frac{\partial z}{\partial x}+\frac{\partial z}{\partial y}=-f' -1+f' =-1\neq z^2"



Conclusion,



"z=-y-f(x-y)\,\,\text{is not a solution to the differential equation}\\[0.3cm]
\frac{\partial z}{\partial x}+\frac{\partial z}{\partial y}=z^2"

QUESTION(iii)


Equation



"\frac{\partial z}{\partial x}\cdot\frac{\partial z}{\partial y}-\left(\frac{\partial z}{\partial y}\right)^2=0"

contains the term "\left(z'_y\right)^2" therefore it is nonlinear.

Conclusion,



"\boxed{\frac{\partial z}{\partial x}\cdot\frac{\partial z}{\partial y}-\left(\frac{\partial z}{\partial y}\right)^2=0-\text{is nonlinear}}"


ANSWER

(i)



"\cos(x+y)p+\sin(x+y)q=z^2-\text{s a quasi-linear equation}"

(ii)



"z=-y-f(x-y)\,\,\text{is not a solution to the differential equation}\\[0.3cm]
\frac{\partial z}{\partial x}+\frac{\partial z}{\partial y}=z^2"

(iii)



"\frac{\partial z}{\partial x}\cdot\frac{\partial z}{\partial y}-\left(\frac{\partial z}{\partial y}\right)^2=0-\text{is nonlinear}"